OpenStudy (anonymous):

for y=4x+2x²-x³ show how to find the value for y when x =2 show your work i have two different answers. one is my brothers one is mine

Directrix (directrix):

What are the two answers that you and the brother got?

OpenStudy (anonymous):

one answer is 8 and the other is 32

Directrix (directrix):

Let's see. If x = 2, then the equation becomes 4*2 +2 * 2² - 2³ =

Directrix (directrix):

4*2 +2 * 2² - 2³ = 8 + 8 - 8 = 8. That is what I got. What do you think?

OpenStudy (anonymous):

i think my brother and everyone else is smarter than me. lol thank you

Directrix (directrix):

I don't think so. Just work the problems in steps. This stuff is not easy for anybody although it may appear that way at times.